OpenStudy (anonymous):

Which terms (linear, nonlinear, quadratic, absolute value) match this equation?

OpenStudy (anonymous):

what equation

OpenStudy (anonymous):

Holdl on.

OpenStudy (anonymous):

\[y = -\left| 3x \right|\] + 1

OpenStudy (anonymous):

absolute value

OpenStudy (anonymous):

\[y = -\left| 3x \right| + 1\]

OpenStudy (anonymous):

thats what the |x| means. it looks like a v shape that the base touches at (0,0)

OpenStudy (anonymous):

when you have a satisfactory answer you can select 'good answer'

OpenStudy (anonymous):

it is also nonlinear

OpenStudy (anonymous):

absolute

OpenStudy (anonymous):

It isn't a linear equation? And sorry, laptop's lagging.

OpenStudy (anonymous):

absolute value is not linear. f(cx)=c*f(x) would be true if it were linear, but it is not for absolute value.

OpenStudy (anonymous):

non parce que il est en bars

OpenStudy (anonymous):

any equatin thay has the absolute value symbol in the equation it will be absolute value. if it didnt have the symbols then yes it wud be linear but in this case its not